# Detecting Edited Knowledge in Language Models

### v2

@@ 5 DEED Baseline Models
To train a detection classifier, we consider different possibilities of available features. Inspired by the probing tasks introduced in Conneau et al. (2018) and the representations inspection by Hernandez et al. (2023), we investigate the use of hidden state representations as features. Additionally, in Kuhn et al. (2023) and Gu et al. (2023), the output probability distributions are used to provide signals for latent semantics. Therefore, we also explore using these distributions as features.
@@ 5.1 Preliminary Analysis
To see how editing affects the hidden state representations, we use Linear Discriminant Analysis (LDA) Duda et al. (2000) to visualize the hidden state representations. Specifically, for each data point (fact) from the test set, we project the hidden state representation of the last token of $p^{\prime}(s,r)$ into a one-dimensional space. We use the representations from the last layer of the respective LLM. We assign the one-dimensional representations random values on the y-axis to reduce point overlap. As shown in Figure 2, the edited and unedited representations are not distinguishable in the unedited model. Editing with ROME and to less extent with MEMIT makes the representations more separable. Editing with MEND and MALMEN keeps the representations linearly inseparable. [^3] We also experimented with PCA and t-SNE, but the representations were still inseparable. This suggests that meta-learning KEs cause subtle changes on representations when editing facts. Similar patterns are observed on all datasets and methods (see Figure 7 in the Appendix).
@@
We investigate the output probability distributions from LLMs when prompting for edited and unedited facts with prompts that contain subjects and relations: $p^{\prime}_{1}(s_{1},r_{1})...p^{\prime}_{n}(s_{n},r_{n})$. We consider a balanced set consisting of the same number of edited and unedited facts. We analyze the probability distribution of each set by calculating the mean of the top 10 probabilities and plotting the kernel density estimate (KDE). We show the KDEs for GPT-J and CounterFact in Figure 3. The figure shows almost identical distributions of low probabilities for both edited and unedited facts in the unedited model NONE (top first row). We notice two peaks for both distributions in NONE: the left one is close to zero, while the right one is around 0.3. As we edit with different methods, we notice that the right peak for edited facts is shifted to the right, while the first one flattens, as the probabilities there become less frequent. The increase in probabilities varies depending on the method; ROME and MEND cause the highest increase, followed by MEMIT and MALMEN, which are used for mass-editing. We also notice that MEND increases the probabilities of unedited facts as well. Editing with MALMEN is the smoothest, since its probability distribution is the closest to that of the unedited facts. This potentially makes facts edited with MALMEN more difficult to detect.
@@
Figure 3: Kernel density estimation for averaged top-10 probabilities.
@@ 5.2 DEED Baselines
Given a fact $(s,r,o,p^{\prime})$, we prompt the given LLM $\mathcal{M}$, which consists of $L$ transformer blocks, with the prompt $p^{\prime}(s,r)$ that consists of $N$ tokens $\{u_{1},\ldots,u_{N}\}$. We use the hidden state representations HS from the last transformer block and the last input token $h_{N,L}=\mathcal{M}[p^{\prime}(s,r)]$ as a first feature for detecting edited facts. As a second feature, we use PD, the top-1000 probabilities $q=(q_{1},\ldots,q_{1000})$ from $\mathcal{M}$ on the same prompt $p^{\prime}(s,r)$. We use the information from the last token, since this is the position where the information about the subject and the relation is condensed Geva et al. (2023).
@@ 5.3 Evaluation
We train an AdaBoost classifier (Freund and Schapire, 1997) with a decision tree as a base classifier using both features, HS and PD, on the DEED datasets (Section 4.2). We show the classification performance of training on 126 instances in the first part of Table 2. The minimum number of edits under all settings was 127 (cf. Table 1). We use half of the edited facts (63) for training, and the rest for testing. Balancing the edited 63 facts with 63 unedited facts yields a training set of 126 examples. As a sanity check, we train a classifier on the same sets of edited and unedited facts using features from the base unedited model. We refer to this classifier as NONE. The classifier NONE shows a fair coin performance on the three metrics, ranging from 46.2 (GPT-J, CounterFact, F1) to 49.8 (GPT2-XL, zsRE, F1). This verifies the correctness of our setup. We report precision, recall, and F1.
@@
We find that the classifier is able to detect edits by locate-and-edit KEs (LE-KEs), i.e., ROME and MEMIT, relatively well, achieving a minimum F1 score of 74.2% (MEMIT, GPT2-XL, zsRE). The highest F1 score of 98% is achieved on (ROME, GPT-J, zsRE). On meta-learning KEs (ML-KEs), i.e., MEND and MALMEN, we observe lower performance in general. On MEND the performance varies between 53.6% (GPT2-XL, CounterFact) and 81.4% (GPT-J, zsRE), whereas on MALMEN the gap between the lowest and highest performance is smaller than the same gap on MEND (62% (GPT2-XL, CounterFact), and 76% (GPT-J, zsRE)). In general, this shows that edits, from the more computationally efficient LE-KEs, are easier to detect. One reason for this discrepancy in performance between these two classes of methods might be that ML-KEs induce changes that generally target the last few layers, while LE-KEs change middle layers. Note for example that MEND increase the probabilities for both edited and unedited facts, while MALMEN slightly increases the probabilities for edited facts (cf. Figure 3). This keeps the distance between edited and unedited facts minimal under ML-KEs, while LE-KEs conduct surgical changes, which increase the probability for edited facts only. This can also be confirmed by the fact that the effects of LE-KEs on unrelated facts are less pronounced compared to ML-KEs (cf. Table 4, Appendix).
Ablation. We conduct an ablation study, where we train only on HS, or PD (second and third part of Table 2 respectively). We notice that using HS leads to low performance on the ML-KEs methods, where the F1 score varies between 48.5% (MALMEN, GPT2-XL, zsRE) and 62.6% (MALMEN, GPT-2XL, CounterFact). On LE-KEs, HS still performs well with a minimum F1 score of 65.3% (MEMIT, GPT-J, CounterFact). On the other hand, detecting edits with PD is more robust and performs well on LE-KEs and ME-KEs. The minimum F1 scores with PD is 65.2% (MALMEN, GPT-J, CounterFact). PD’s robustness might be explained by the fact that it is derived from a late position in the LLM, namely the language modeling head, which is responsible for the final outputs. Another important observation is that using PD individually can lead to better performance than combining it with HS in some cases (e.g., the F1 score on (MEND, GPT2-XL, CounterFact) is 53.6% when using both features, but 73% when using only PD). This observation suggests that PD and HS could be combined in better ways. We leave exploring the interplay between HS and PD to future work.
@@
Figure 4: Detection performance with PD on CounterFact (left) and zsRE (right) when varying training set size.
@@
Figure 4 shows the performance of different training set sizes with PD. Exhaustive results are presented in Table 7 in the Appendix. In general, the classification performance starts reaching a plateau after 50 training instances. For example, on zsRE and MEMIT-edited knowledge on GPT-J (the green line in the right figure), F1 reaches 85.4% after 50 training instances, while the highest F1 of 87.2% is achieved after 300 training data points. Therefore, we highlight that this classifier remains effective even when the training data consists of less than one hundred examples. We also observe that facts that are edited with ML-KEs are not only harder to detect than facts edited with LE-KEs (as noted before), but also that scaling the size of the training set does not improve performance.
HS from different layers.
To investigate the effect of HS from different layers on the classification performance, we train a simple logistic regression classifier with L1 regularization on HS from different layers. We consider only locate-and-edit methods (ROME and MEMIT) for this experiment, since HS performs well on this category of methods. We show the F1 values based on hidden states representations from different layers in Figure 5. First, for both models, editing methods, and datasets, the representations start becoming separable, i.e., the classification performance is above random guess, after the edited layers or shortly thereafter. For example, the F1 for detecting ROME edits starts improving from the layer 7-8 in GPT-J and the layer 19 in GPT2-XL (ROME changes layer 7 in GPT-J, and layer 19 in GPT2-XL). We observe the same effect with MEMIT (MEMIT changes the layers 5-10 in GPT-J and the layers 15-19 in GPT2-XL). Even though the representations from the last layer perform well in detecting edits, we notice that the F1 reaches its peak at preceding layers. For ROME, the highest performance can be seen at training on representations from the layer 13-27 (GPT-J) or from the layers 23-47 (GPT2-XL). For MEMIT, the highest F1 is reached with representations from layer 17-27 (GPT-J), and layer 28-47 (GPT2-XL). This suggests that utilizing outputs from the last layer for detecting edited knowledge might not always be the optimal choice. Further, we observe that using the penultimate layer’s representations is more effective than using the last layer’s to detect edited knowledge. For example, for edited GPT-J on CounterFact (first subplot in Figure 5), representations from the penultimate layer achieve an F1 of 90% and 72% for detecting ROME-edited and MEMIT-edited knowledge, respectively, higher than 85% (ROME) and 67% (MEMIT) when using the representation from the last layer. LE-KEs might be causing the attribute (object) extraction phase Geva et al. (2023) that is responsible for extracting the object to happen early in the model. This early extraction is often associated with high probabilities (cf. Section 5.1), causing edits from LE-KEs to be easier to detect. In general, this experiment shows that the effect of edits from LE-KEs starts becoming visible from the changed layers, increases strongly thereafter, and decreases slightly at the last layer. This demonstrates that using representations from the last layer is not the optimal choice; utilizing representations from other layers can further improve performance.
6 Detection in Challenging Settings
6.1 Cross-Domain DEED
Obtaining a training set for the detection classifier requires knowing instances of edited facts in the edited model. However, this assumption may not always hold. Consider a scenario where an open-source model is further fine-tuned for a specific task of interest to the user, yet no explicit information is provided regarding the edited facts. Many models are fine-tuned with instructions or reinforcement learning from human feedback (RLHF), and are made publicly available Lai et al. (2023). In these cases, we have access to the original model (from which we could obtain our training set), but do not know which facts have been edited into the fine-tuned model. To address such cases, we assess the performance of the classifier, which uses HS and PD, in a cross-domain setting, where the classifier is trained on representations from one edited model and then applied to detecting edited knowledge from a separate edited model (which is a fine-tuned version of the original model). The CD edited models have the same architecture, but different weights. To simulate this scenario, we use representations from edited GPT-J and GPT2-XL to train the classifier and test it in detecting edited knowledge from fine-tuned versions of GPT-J [^4] https://huggingface.co/togethercomputer/GPT-JT-6B-v1 and GPT2-XL [^5] https://huggingface.co/lgaalves/gpt2-xl_lima. Note that the facts in the training and test sets are still disjoint. For this experiment, we consider ROME, MEMIT and MALMEN.
@@ 6.1 Cross-Domain DEED
As shown in Figure 6, the performance on cross-domain edits (CD) is on par with performance on in-domain data (ID). For example, on zsRE, CD F1 on ROME-edited knowledge on GPT2-XL lima is 95%, similar to the ID F1 of 96.6%. Full results are shown in Table 6 in the Appendix. This makes our baseline for detecting edited knowledge highly practical, as it can be leveraged to detect edits in fine-tuned models, while making use of training data that is derived from their counterpart original model.
6.2 Robustness Evaluation
Generator zsRE CounterFact
Model Editor Pr. Rec. F1 Pr. Rec. F1
default MEMIT 87.7 89.8 88.7 79.0 76.7 77.9
MALMEN 76.7 75.3 76.0 62.0 62.0 62.0
same object MEMIT 58.5 91.1 71.3 72.7 72.9 72.8
MALMEN 57.5 57.9 57.7 56.4 63.6 59.8
Table 3: Classification performance on detecting edited knowledge for GPT-J. Facts in the training and test data end with the same object. Classification is done based on HS and PD.
@@ 7 Conclusion
In this paper, we introduced a novel task: detecting edited knowledge in language models. We see promising applications in uncovering potential malicious model modification and enhancing the transparency of LLMs’ generations. We instantiated the task with several baselines, which use hidden state representations and probability distributions as input features. These baselines demonstrate a straightforward yet effective approach to the task. Our evaluation across two LLMs, four editing techniques (2 LE-KEs and 2 ML-KEs), and two datasets, shows that the classifiers can reliably detect edits, even in scenarios characterized by limited training data. Moreover, we showed that these baselines generalize to cross-domain settings, highlighting their practical utility, and that distinguishing between edited and unedited facts with the same object is more challenging, suggesting potential future work directions. In future work, we aim to explore detecting if an LLM has undergone any edits at all, study generalization to other KEs and more constrained settings, where we only have access to the LLM’s outputs.
@@ 8 Limitations
In our current setting, we assume we know the editing method. However, the editing method might not be known and an edited model can be modified by multiple editing methods. In future work, we will investigate the generalization of our baseline classifiers between different KEs.
@@ Appendix A Complementary Results
In Table 5, we report the number of training and test instances used in our default setting. We keep the number of training instances constant across all settings for consistency. Results for varying sizes of the training set are presented in Table 7. The result for the CD evaluation is shown in Table 6. We show some prompt examples that are used for editing and evaluating edits in Table 8. We show the HS representations in a low-dimensional space in Figure 7.
@@
Generator zsRE CounterFact
Model Editor Acc. Pr. Rec. F1 Acc. Pr. Rec. F1
GPT-JT ROME 90.7 87.1 95.5 91.1 92.4 90.9 94.6 92.5
MEMIT 80.5 84.3 75.0 79.4 78.4 76.6 81.8 79.1
MALMEN 77.2 80.7 71.7 75.9 67.6 65.6 74.1 69.6
GPT2-XL lima ROME 94.8 91.4 99.0 95.0 91.5 89.6 93.8 91.6
MEMIT 74.0 79.3 65.0 71.4 77.2 77.3 77.1 77.2
MALMEN 74.8 73.4 77.70 75.5 70.4 68.5 75.5 71.8
Table 6: CD (Cross-Domain) classification performance. Training data is from GPT2-XL and GPT-J, whereas test data is from GPT2-XL lima and GPT-JT respectively.
@@
Table 7: DEED (detecting edited knowledge) performance with increasing amount of training data.

### v3

@@ 5 DEED Baseline Models
To train a detection classifier, we consider different options of available features. Inspired by the probing tasks introduced in Conneau et al. (2018) and the representations inspection by Hernandez et al. (2023), we investigate the use of hidden state representations as features. Additionally, in Kuhn et al. (2023) and Gu et al. (2023), the output probability distributions are used to provide signals for latent semantics. Therefore, we also explore using these distributions as features.
@@ 5.1 Preliminary Analysis
To see how editing affects the hidden state representations, we use Linear Discriminant Analysis (LDA) Duda et al. (2000) to visualize the hidden state representations. Specifically, for each data point (fact) from the test set, we project the hidden state representation of the last token of $p^{\prime}(s,r)$ into a one-dimensional space. We use the representations from the last layer of the respective LLM. We assign the one-dimensional representations random values on the y-axis to reduce point overlap. As shown in Figure 2, the edited and unedited representations are not distinguishable in the unedited model. Editing with ROME and to less extent with MEMIT makes the representations more separable. We attribute this separability to the high confidence of LLMs in the edited facts, as shown later in Section 5.3. Editing with MEND and MALMEN keeps the representations linearly inseparable. [^4] We also experimented with PCA and t-SNE, but the representations were still inseparable. This suggests that meta-learning KEs cause subtle changes on representations when editing facts. Similar patterns are observed on all datasets and methods (see Figure 6 in the Appendix).
@@
We investigate the output probability distributions for the next token from LLMs when prompting for edited and unedited facts with prompts that contain subjects and relations: $p^{\prime}_{1}(s_{1},r_{1})...p^{\prime}_{n}(s_{n},r_{n})$. We consider a balanced set consisting of the same number of edited and unedited facts. We analyze the probability distribution of each set by calculating the mean of the top 10 probabilities and plotting the kernel density estimate (KDE). We show the KDEs for GPT-J and CounterFact in Figure 3. The figure shows almost identical distributions of low probabilities for both edited and unedited facts in the unedited model NONE. We notice two peaks for both distributions in NONE: the left one is close to zero, while the right one is around 0.3. As we edit with different methods, we notice that the right peak for edited facts is shifted to the right, while the first one flattens, as the probabilities there become less frequent. The increase in probabilities varies depending on the method; ROME and MEND cause the highest increase, followed by MEMIT and MALMEN, which are used for mass-editing. We also notice that MEND increases the probabilities of unedited facts as well. Editing with MALMEN is the smoothest, since its probability distribution is the closest to that of the unedited facts. This potentially makes facts edited with MALMEN more difficult to detect.
@@
Figure 3: Kernel density estimation for averaged top-10 probabilities for the next token (GPT-J, CounterFact).
@@ 5.2 DEED Baselines
Given a fact $(s,r,o,p^{\prime})$, we prompt the given LLM $\mathcal{M}$, which consists of $L$ transformer blocks, with the prompt $p^{\prime}(s,r)$ that consists of $N$ tokens $\{u_{1},\ldots,u_{N}\}$. We use the hidden state representations HS from the last transformer block and the last input token $h_{N,L}=\mathcal{M}[p^{\prime}(s,r)]$ as a first feature for detecting edited facts. We use the information from the last token, since this is the position where the information about the subject and the relation is condensed Geva et al. (2023). As a second feature, we use PD, the top-1000 probabilities of the next token $q=(q_{1},\ldots,q_{1000})$ from $\mathcal{M}$ on the same prompt $p^{\prime}(s,r)$.
@@ 5.3 Evaluation
We train an AdaBoost classifier (Freund and Schapire, 1997) with a decision tree as a base classifier using both features, HS and PD, on the DEED datasets (Section 4.2). We show the classification performance of training on 126 instances in the first part of Table 2. The minimum number of edits under all settings was 127 (cf. Table 1). We use half of the edited facts (63) for training, and the rest for testing. Balancing the edited 63 facts with 63 unedited facts yields a training set of 126 examples. We keep the size of the training set constant across all settings for fair comparison, and experiment later with smaller and larger training sets (cf. Figure 4). As a sanity check, we train a classifier on the same sets of edited and unedited facts using features from the base unedited model. We refer to this classifier as NONE. The classifier NONE shows a fair coin performance on the three metrics, ranging from 46.2 (GPT-J, CounterFact, F1) to 49.8 (GPT2-XL, zsRE, F1). This verifies the correctness of our setup. We report precision, recall, and F1.
@@
We find that the classifier is able to detect edits by locate-and-edit KEs (LE-KEs), i.e., ROME and MEMIT, relatively well, achieving a minimum F1 score of 74.2% (MEMIT, GPT2-XL, zsRE). The highest F1 score of 98% is achieved on (ROME, GPT-J, zsRE). On meta-learning KEs (ML-KEs), i.e., MEND and MALMEN, we observe lower performance in general. On MEND the performance varies between 53.6% (GPT2-XL, CounterFact) and 81.4% (GPT-J, zsRE), whereas on MALMEN the gap between the lowest and highest performance is smaller than the same gap on MEND (62% (GPT2-XL, CounterFact), and 76% (GPT-J, zsRE)). In general, this shows that edits, from the more computationally efficient LE-KEs, are easier to detect. One reason for this discrepancy in performance between these two classes of methods might be that ML-KEs induce changes that generally target the last few layers, while LE-KEs change middle layers. Note for example that MEND increases the probabilities for both edited and unedited facts, while MALMEN slightly increases the probabilities for edited facts (cf. Figure 3). This keeps the distance between edited and unedited facts minimal under ML-KEs, while LE-KEs conduct surgical changes that increase the probability for edited facts only. This can also be confirmed by the fact that the effects of LE-KEs on unrelated facts are less pronounced compared to ML-KEs (cf. Table 4, Appendix).
Ablation. We conduct an ablation study, where we train only on HS, or PD (second and third part of Table 2 respectively). We notice that using HS leads to low performance on the ML-KEs methods, where the F1 score varies between 48.5% (MALMEN, GPT2-XL, zsRE) and 62.6% (MALMEN, GPT-2XL, CounterFact). On LE-KEs, HS still performs well with a minimum F1 score of 65.3% (MEMIT, GPT-J, CounterFact). In contrast, detecting edits with PD is more robust and performs well on LE-KEs and ML-KEs. The minimum F1 scores with PD is 65.2% (MALMEN, GPT-J, CounterFact). PD’s robustness might be explained by the fact that it is derived from a late position in the LLM, namely the language modeling head, which is responsible for the final outputs. Another important observation is that using PD individually can lead to better performance than combining it with HS in some cases (e.g., the F1 score on (MEND, GPT2-XL, CounterFact) is 53.6% when using both features, but 73% when using only PD). This observation suggests that PD and HS could be combined in better ways. We leave exploring the interplay between HS and PD to future work. We further explore using HS from different layers in Appendix A.1, observing that the representations start becoming separable after the edited layers, and that the representations from some layers (e.g., the penultimate layer) perform better than the representations of the last layer.
@@
Figure 4: Detection performance with PD on zsRE when varying the training set size.
@@
Figure 4 shows the performance of different training set sizes with PD. Exhaustive results are presented in Table 7 in the Appendix. In general, the classification performance starts reaching a plateau after 50 training instances. For example, on MEMIT-edited facts in GPT-J (green circles), F1 reaches 85.4% after 50 training instances, while the highest F1 of 87.2% is achieved after 300 training data points. Therefore, we emphasize that this classifier remains effective even when the training data consists of less than one hundred examples. We also observe that facts that are edited with ML-KEs are not only harder to detect than facts edited with LE-KEs (as noted before), but also that scaling the size of the training set does not improve performance.
6 Cross-Domain DEED
@@
Obtaining a training set for the detection classifier requires knowing instances of edited facts in the edited model. However, this assumption may not always hold. Consider a scenario where an open-source model is further fine-tuned for a specific task of interest to the user, yet no explicit information is provided regarding the edited facts. Many models are fine-tuned with instructions or reinforcement learning from human feedback (RLHF), and are made publicly available Lai et al. (2023). In these cases, we have access to the original model (from which we could obtain our training set), but do not know which facts have been edited into the fine-tuned model. To address such cases, we assess the performance of the classifier, which uses HS and PD, in a cross-domain setting, where the classifier is trained on representations from one edited model and then applied to detecting knowledge edits from a separate edited model (which is a fine-tuned version of the original model). The CD edited models have the same architecture, but different weights. To simulate this scenario, we use representations from edited GPT-J and GPT2-XL to train the classifier and test it in detecting knowledge edits from instruction-tuned versions of GPT-J [^5] https://huggingface.co/togethercomputer/GPT-JT-6B-v1 and GPT2-XL. [^6] https://huggingface.co/lgaalves/gpt2-xl_lima Note that the facts in the training and test sets are still disjoint. For this experiment, we consider ROME, MEMIT and MALMEN.
As shown in Figure 5, the performance on cross-domain edits (CD) is on par with performance on in-domain data (ID). For example, on zsRE, CD F1 on ROME-edited facts on GPT2-XL lima is 95%, similar to the ID F1 of 96.6%. Full results with HS/PD are shown in Table 6 in the Appendix. This makes our DEED baseline highly practical, as it can be leveraged to detect edits in fine-tuned models, while making use of training data that is derived from their counterpart original models.
Cross-dataset generalization.
The results of our previous experiment indicate that detection classifiers trained on edits in one model generalize to edits in other models of the same family. To further explore out-of-distribution scenarios, we impose an additional constraint by examining whether detection classifiers trained on edits from one dataset (e.g., CounterFact) generalize to edits from another dataset (e.g., zsRE). Here, we maintain the assumption from our previous experiment that the base LLM is known, but introduce a new constraint: the dataset used to conduct edits on the test LLM is unknown, necessitating training on one dataset and testing on another. The primary objective of this experiment is to investigate whether edit detection classifiers generalize effectively across datasets.
As shown in Table 3 the detection classifiers generalize well to unseen datasets in most cases. One exception is generalization from zsRE to CounterFact with MALMEN, where we see a dramatic decrease in performance. One reason for this might be the smooth changes induced by MALMEN (cf. Section 5.1). These changes are dataset-specific, which makes generalization more difficult (cf. Figure 9(b) and 9(c) in the Appendix).
Method zsRE CounterFact
GPT-J ROME 82.7 89.6
MEMIT 73.4 76.8
MALMEN 63.7 25.9
GPT2-XL ROME 84.5 90.8
MEMIT 70.4 72.7
MALMEN 62.1 6.8
Table 3: F1 scores in the cross-datasets setting. Classifiers are trained on HS and PD from one dataset and tested on another dataset in CD setting. Tests sets are shown in the header.
@@ 7 Conclusion
In this paper, we introduced a novel task: detecting knowledge edits in language models. We see promising applications in uncovering potential malicious model modification and enhancing the transparency of LLMs’ generations. We instantiated the task with several baselines, which use hidden state representations and probability distributions as input features. These baselines demonstrate a straightforward yet effective approach to the task. Our evaluation across two LLMs, four editing techniques (2 LE-KEs and 2 ML-KEs), and two datasets, shows that the classifiers can reliably detect edits, even in scenarios characterized by limited training data. Moreover, we showed that these baselines generalize to cross-domain settings, where training and test data come from different models and datasets, highlighting their practical utility. Our work gives rise to several future work directions: 1) detecting if an LLM has undergone any edits at all; 2) studying detection generalization across KEs; 3) investigating more constrained settings, where one has only access to the LLM’s outputs; 4) combating malicious edits.
@@ 8 Limitations
In our current setting, we assume we know the editing method. However, the editing method might not be known and an edited model can be modified by multiple editing methods. In future work, we will investigate the generalization of our baseline classifiers between different KEs, and detecting edits in models that have edited with several KEs.
@@
Acknowledgment
This research was partially funded by the German Federal Ministry of Education and Research (BMBF) as German Academic Exchange Service (DAAD) project DEED under Grant No. 30001797. This work was conducted while Zhixue Zhao was an AInet Fellow, funded under the DAAD Fellowship Program (No. 57733361).
@@ Appendix A Complementary Results
In Table 5, we report the number of training and test instances used in our default setting. We keep the number of training instances constant across all settings for consistency. Results for varying sizes of the training set are presented in Table 7 and Figure 8 (CounterFact). The result for the CD evaluation is shown in Table 6. We show some prompt examples that are used for editing and evaluating edits in Table 10. We show the HS representations in a low-dimensional space in Figure 6. We show KDE visualizations of PD across all datasets, models, and KEs in Figure 9.
@@
zsRE CounterFact zsRE CounterFact zsRE CounterFact
KE Pr. Rec. F1 Pr. Rec. F1 Pr. Rec. F1 Pr. Rec. F1 Pr. Rec. F1 Pr. Rec. F1
GPT-JT NONE 49.2 27.8 35.6 50.4 49.5 50.0 48.4 32.8 39.1 49.6 43.5 46.4 0.0 0.0 0.0 50.2 39.7 44.3
ROME 87.1 95.5 91.1 90.9 94.2 92.5 99.8 74.2 85.1 91.6 87.8 89.6 84.8 84.5 84.7 89.3 94.0 91.6
MEMIT 84.3 75.0 79.4 76.6 81.8 79.1 80.5 56.2 66.2 69.9 76.5 73.0 91.6 68.7 78.5 79.3 84.7 81.9
MALMEN 80.7 71.6 75.9 65.6 74.1 69.6 50.4 34.6 41.1 57.8 69.1 63.0 72.8 88.4 79.9 82.4 54.0 65.2
GPT2 lima NONE 49.0 56.1 52.3 51.5 45.9 48.5 50.1 41.8 45.6 49.4 42.0 45.4 50.6 93.0 65.6 51.2 39.2 44.4
ROME 91.4 99.0 95.0 89.6 93.8 91.6 97.8 93.2 95.4 94.5 83.0 88.4 63.6 86.9 73.4 88.7 91.3 90.0
MEMIT 79.2 65.0 71.4 77.3 77.1 77.2 76.9 61.9 68.6 73.7 60.9 66.7 74.0 73.2 73.6 76.9 76.4 76.7
MALMEN 73.4 77.7 75.5 68.5 75.5 71.8 48.3 53.9 51.0 65.0 53.1 58.4 73.6 92.0 81.8 68.1 65.3 66.7
Table 6: CD (Cross-Domain) classification performance. Training data is from GPT2-XL and GPT-J, whereas test data is from GPT2-XL lima (GPT2 lima) and GPT-JT respectively.
@@
Table 7: DEED (detecting knowledge edits) performance with increasing amount of training data.
@@
A.1 HS from Different Layers
To investigate the effect of HS from different layers on the classification performance, we train a simple logistic regression classifier with L1 regularization on HS from different layers. We consider only locate-and-edit methods (ROME and MEMIT) for this experiment, since HS performs well on this category of methods. We show the F1 values based on hidden states representations from different layers in Figure 7. First, for both models, editing methods, and datasets, the representations start becoming separable, i.e., the classification performance is above random guess, after the edited layers or shortly thereafter. For example, the F1 for detecting ROME edits starts improving from the layer 7-8 in GPT-J and the layer 19 in GPT2-XL (ROME changes layer 7 in GPT-J, and layer 19 in GPT2-XL). We observe the same effect with MEMIT (MEMIT changes the layers 5-10 in GPT-J and the layers 15-19 in GPT2-XL). Even though the representations from the last layer perform well in detecting edits, we notice that the F1 reaches its peak at preceding layers. For ROME, the highest performance can be seen at training on representations from the layer 13-27 (GPT-J) or from the layers 23-47 (GPT2-XL). For MEMIT, the highest F1 is reached with representations from layer 17-27 (GPT-J), and layer 28-47 (GPT2-XL). This suggests that utilizing outputs from the last layer for detecting knowledge edits might not always be the optimal choice. Further, we observe that using the penultimate layer’s representations is more effective than using the last layer’s to detect knowledge edits. For example, for edited GPT-J on CounterFact (first subplot in Figure 7), representations from the penultimate layer achieve an F1 of 90% and 72% for detecting ROME-edited and MEMIT-edited knowledge, respectively, higher than 85% (ROME) and 67% (MEMIT) when using the representation from the last layer. LE-KEs might be causing the attribute (object) extraction phase Geva et al. (2023) that is responsible for extracting the object to happen early in the model. This early extraction is often associated with high probabilities (cf. Section 5.1), causing edits from LE-KEs to be easier to detect. In general, this experiment shows that the effect of edits from LE-KEs starts becoming visible from the changed layers, increases strongly thereafter, and decreases slightly at the last layer. This demonstrates that using representations from the last layer is not the optimal choice; utilizing representations from other layers can further improve performance.
A.2 Robustness Evaluation
Generator zsRE CounterFact
Model Editor Pr. Rec. F1 Pr. Rec. F1
default MEMIT 87.7 89.8 88.7 79.0 76.7 77.9
MALMEN 76.7 75.3 76.0 62.0 62.0 62.0
same object MEMIT 58.5 91.1 71.3 72.7 72.9 72.8
MALMEN 57.5 57.9 57.7 56.4 63.6 59.8
Table 8: Classification performance on detecting knowledge edits for GPT-J. Facts in the training and test data end with the same object. Classification is done based on HS and PD.
Figure 8: Detection performance with PD on CounterFact when varying training set size.
(a) zsRE and GPT-J.
(b) CounterFact and GPT2-XL.
(c) zsRE and GPT2-XL.
Figure 9: KDE of edited and non-edited facts.
A.3 Detection by Top 10 Output Probabilities
We experiment with using only the top 10 output probabilities instead of the top 1,000 output probabilities. The differences in performance (top 10 performance - top 1000 performance)) in Table 9 show that the top-10 output probabilities works well with minor degradation in performance in most cases. This means that our baselines can also be used on proprietary LLMs that output the top 10 output probabilities.
Generator zsRE CounterFact
Model Editor Pr. Rec. F1 Pr. Rec. F1
GPT-J NONE -2.6 6.7 2.4 -2.4 -4.5 -3.5
ROME -0.4 -0.4 -0.4 1.5 -1.5 0.0
MEMIT -0.6 -0.9 -0.7 -1.7 2.6 0.6
MEND -0.2 3.2 1.3 0.6 -0.4 0.3
MALMEN 2.6 -1.9 0.5 4.0 -0.5 1.6
GPT2-XL NONE -0.1 0.3 0.1 -0.8 0.0 -0.4
ROME 0.1 -2.7 -1.3 2.0 -5.2 -1.5
MEMIT -1.4 -5.7 -3.8 -7.0 -1.7 -4.4
MEND -1.0 -5.8 -2.9 0.1 0.1 0.1
MALMEN 0.0 9.2 4.7 -4.6 -14.1 -9.4
Table 9: Differences in classification performance for detecting knowledge edits: top 10 probabilities minus top 1,000 probabilities